Food Food A (Y) X 200 Y-200 X-6667 Y=100 dles Oil(x) United States United States Canada S. The production possibility frontiers of two counties are represented on the graphs above. The production possibility frontiers are represented as straight lines. Let Good X be units of oil and Good Y be units of food. The United States is currently producing and consuming the goods combination represented by Point A. Canada is currently producing and consuming the goods combination represented by Point B. a. Does either country have an absolute advantage in production? Explain. b. Which country has a comparative advantage in the production of oil (Good x)? Which has a comparative advantage in the production of food (Good Y)? Explain and show your work. Would the two countries find it mutually beneficial if the United States were to trade 25 units of oil (good X) to Canada in return for Canada sending 50 units of food (good Y) to the United States? Explain. c.
Answer #1

a) Absolute Advantage:

Absolute advantage is the ability of a country to produce a good or service at a lower per unit cost as compared to any other country that produces same good or service.

United States has an absolute advantage in the production of Oil and Food.

b) A country has a comparative advantage in producing that good if the opportunity cost of producing that good is lower in that country as compared to another country.

United States;

400 units of Oil = 400 units of Food

1 unit of Oil = 400/400 unit = 1 unit of Food

1 unit of Food = 400/400 = 1 unit of Oil

Opportunity cost of producing 1 unit of Oil is 1 unit of Food and 1 unit of Food is 1 unit of Oil.

Canada;

100 units of Oil = 300 units of Food

1 unit of Oil = 300/100 unit = 3 unit of Food

1 unit of Food = 100/300 = 0.33 unit of Oil

Opportunity cost of producing 1 unit of Oil is 3 unit of Food and 1 unit of Food is 0.33 unit of Oil.

United States has comparative advantage in the production of Oil while Canada has comparative advantage in the production of Food.

With Specialization; United States will only produce Oil and maximum production = 400 units

Canada will produce only Food and production = 300

c) Trade; 25 units of Oil = 50 units of Food

United States; Oil = 400 - 25 = 375 units And Food = 50 units

Canada; Oil = 25 units And Food = 300 - 50 = 250 units

Food Food A (Y) X200 Y-200 X-6667 Y-100 oil (X) Oil(x) United States

Both countries are able to consume good which lies outside PPF because of trade. So, both countries are better off.